OpenStudy (anonymous):

I'm really confused on what the limit would be here:

OpenStudy (anonymous):

lim (1)/(x-10)^2 x-->10

OpenStudy (anonymous):

I think there's no limit but wolfram says it's -1/90 how is that possible?

OpenStudy (luigi0210):

Do you accept \(\pm \infty/DNE\)?

OpenStudy (anonymous):

What do you mean do I accept those? Yea I've dealt with them before if that's what you mean

OpenStudy (anonymous):

ohhh I see my mistake but I have to explain why it is infinity

OpenStudy (phi):

as x->10 x-10 approaches 0 when you square it, you get +0 from either direction divide by a number that approaches 0 gives you a number that tends to infinity

OpenStudy (anonymous):

but isn't 1/0 undefined. that's what confuses me

OpenStudy (luigi0210):

If you wanna be exact.. \(\lim_{x \rightarrow 10^-} f(x)=\infty \) and \(\lim_{x \rightarrow 10^+} f(x)=\infty \), \(\lim_{x \rightarrow 10^-} f(x)=\lim_{x \rightarrow 10^+} f(x)\) So the limit is \(\infty\) :P

OpenStudy (anonymous):

but what makes it infinity like what do you plug in to the equation and how do you expand it

OpenStudy (anonymous):

okay thanks :)

OpenStudy (phi):

The idea is we do *not* divide by 0. rather, we ask what does the quotient *approach* as x approaches 0.

OpenStudy (anonymous):

yea i understood thank you so much